# Perturbative Chern-Simons invariants from non-acyclic flat connections

### v1

@@ 1. Introduction
In his pioneering work [13], Witten was studying the path integral of Chern-Simons theory
@@ 2. The construction of the invariant
Let $M$ be a closed oriented smooth 3-manifold. Fix $G$ a compact simply-connected Lie group with Lie algebra $\mathfrak{g}$. Also fix a Riemannian metric $g$ on $M$ (this is auxiliary data: the invariants we construct will turn out to be independent of $g$).
@@
(5) $\Phi_{\Gamma}=\int_{C_{V}(M)}\Big\langle\bigwedge_{\mathrm{edges\;}e=(uv)}\pi_{uv}^{*}\widehat{\eta}\;\wedge\;\bigwedge_{\mathrm{short\;loops\;}e=(vv)}\pi^{*}_{v}\widehat{\eta}^{\Delta}\;\wedge\;\bigwedge_{\mathrm{leaves\;}l}\pi^{*}_{v(l)}\widehat{i}(\mathsf{a}^{2}),\bigotimes_{\mathrm{vertices}}f\Big\rangle_{\Gamma}.$
@@
  • $\langle-,-\rangle_{\Gamma}$ is the extension of the Killing form on $\mathfrak{g}$ to the pairing between $\mathfrak{g}^{\otimes N}$ and $\mathfrak{g}^{\otimes N}$ with $N$ the number of half-edges in $\Gamma$, according to the incidence of half-edges to vertices in $\Gamma$.
@@
$\Theta=\frac{1}{2\epsilon}\int_{M}\langle\widehat{i}(\mathsf{a}^{2})\stackrel{{\scriptstyle\wedge}}{{,}}d\widehat{i}(\mathsf{a}^{2})\rangle=\frac{\epsilon}{2}\langle\mathsf{a}^{2},p\,\mathrm{ad}^{*}_{i(\zeta)}GdG\mathrm{ad}^{*}_{i(\zeta)}i(\mathsf{a}^{2})\rangle,$
@@
We define the “Chern-Simons volume form” on $\mathcal{M}^{\prime}$ as the Berezin integral
(8) $Y=\tau_{A}^{\frac{1}{2}}\int_{H_{A}^{1}[1]\oplus H_{A}^{2}[-1]}\mathcal{D}\zeta\,\mathcal{D}\mathsf{a}^{2}\;e^{\langle\zeta,\mathsf{a}^{2}\rangle+W-W_{0}}\quad\in\Omega^{\mathrm{top}}(\mathcal{M}^{\prime})[[\epsilon]].$
@@
be the Chern-Simons functional evaluated on the Levi-Civita connection associated to the metric $g$, trivialized using $\phi$.
@@
(9) $c(\epsilon)=\frac{h^{\vee}\dim\mathfrak{g}}{24(2\pi)^{2}}\epsilon+O(\epsilon^{3})\;\in\epsilon\mathbb{R}[[\epsilon^{2}]]$
@@
does not depend on the metric $g$ and thus is an invariant of the framed oriented 3-manifold $(M,\phi)$. In (9), $h^{\vee}$ is the dual Coxeter number of $\mathfrak{g}$.
@@ 3. The lowest (nontrivial) degree invariant
(13) $\Lambda_{1}=\int_{\mathcal{M}^{\prime}_{\alpha}}\tau_{A}^{\frac{1}{2}}\left(\Phi_{1}+\Phi_{2}+\Phi_{3}+\Phi_{4}+\Phi_{5}+\Phi_{6}+\frac{h^{\vee}\dim\mathfrak{g}}{24(2\pi)^{2}}S_{\mathrm{grav}}(g,\phi)\right)$
@@
$\Phi_{1}$ $=$ $\frac{1}{12}\int_{C_{2}(M)}\langle\eta\wedge\eta\wedge\eta,f\otimes f\rangle_{\Gamma_{1}},$
$\Phi_{2}$ $=$ $\frac{1}{8}\int_{C_{2}(M)}\langle\pi_{1}^{*}\eta^{\Delta}\wedge\eta\wedge\pi_{2}^{*}\eta^{\Delta},f\otimes f\rangle_{\Gamma_{2}},$
$\Phi_{3}$ $=$ $\sum_{a}\frac{1}{2}\int_{M}\langle[K\mathrm{ad}^{*}_{\chi_{a}}G]^{\Delta}\wedge\chi^{a},f\rangle,$
$\Phi_{4}$ $=$ $\sum_{a}\frac{1}{2}\int_{M}\langle\eta^{\Delta}\wedge G\mathrm{ad}^{*}_{\chi_{a}}\chi^{a},f\rangle,$
$\Phi_{5}$ $=$ $\sum_{a,b}\frac{1}{2}\int_{M}\langle\chi^{a}\stackrel{{\scriptstyle\wedge}}{{,}}\mathrm{ad}^{*}_{\chi_{b}}GdG\mathrm{ad}^{*}_{\chi_{a}}\chi^{b}\rangle,$
$\Phi_{6}$ $=$ $\sum_{a,b}\frac{1}{2}\int_{M}\langle\chi^{a}\stackrel{{\scriptstyle\wedge}}{{,}}\mathrm{ad}^{*}_{\chi_{a}}GdG\mathrm{ad}^{*}_{\chi_{b}}\chi^{b}\rangle.$
Here for each point of $\mathcal{M}^{\prime}_{\alpha}$ we choose arbitrarily a basis $\chi_{a}$ in $A$-harmonic 1-forms, and $\chi^{a}$ is the dual basis of harmonic 2-forms. The object $K\mathrm{ad}^{*}_{\chi_{a}}G]^{\Delta}\in\Omega^{1}(M,\mathfrak{g}\otimes\mathfrak{g})$ is the (regularized) evaluation of the integral kernel of the operator in brackets on the diagonal. [^9] One can also write $\Phi_{3}$ as an integral over the two-point configuration space $\sum_{a}\frac{1}{2}\int_{C_{2}(M)}\langle\eta\wedge[\mathrm{ad}^{*}_{\chi_{a}}G]\wedge\pi_{2}^{*}\chi^{a},\varkappa\otimes f\rangle$}, with $\cdots]\in\Omega^{2}(C_{2}(M),\mathfrak{g}\otimes\mathfrak{g})$ standing for the integral kernel of the operator in brackets and $\varkappa\in\mathfrak{g}\otimes\mathfrak{g}$ the canonical element corresponding to the Killing form on $\mathfrak{g}$. Yet another way to write $\Phi_{3}$ is as a regularized supertrace $\sum_{a}\frac{1}{2}\mathrm{Str}_{\Omega^{\bullet}(M,\mathfrak{g})}^{\mathrm{reg}}\left(K\mathrm{ad}^{*}_{\chi_{a}}G\mathrm{ad}_{\chi^{a}}\right)$. Expressions $\Phi_{1},\ldots,\Phi_{6}$ correspond to the six graphs in Figure 1. Graphs $\Gamma_{1},\Gamma_{2}$ are as in the usual Axelrod-Singer two-loop Chern-Simons invariant (at an acyclic flat connection), whereas the remaining graphs $\Gamma_{3},\ldots,\Gamma_{6}$ account (in an appropriate sense) for the effect of globalization on the moduli space $\mathcal{M}^{\prime}$.
@@
(14) $\Xi=\sum_{a}\iota_{\chi_{a}}\tau_{A}^{\frac{1}{2}}\cdot\Big(\int_{M}\frac{1}{2}\langle\eta^{\Delta}\wedge K\lambda_{\delta g}\chi^{a},f\rangle+\int_{M}\frac{1}{2}\langle[K\lambda_{\delta g}K]^{\Delta}\wedge\chi^{a},f\rangle\\ +\sum_{b}\int_{M}\langle\chi^{b}\stackrel{{\scriptstyle\wedge}}{{,}}\lambda_{\delta g}G\mathrm{ad}^{*}_{\chi_{b}}\chi^{a}\rangle+\sum_{b}\int_{M}\langle\chi^{b}\stackrel{{\scriptstyle\wedge}}{{,}}\mathrm{ad}^{*}_{\chi_{b}}G\lambda_{\delta g}\chi^{a}\rangle\Big)\quad\in\Omega^{\mathrm{top}-1}(\mathcal{M}^{\prime}_{\alpha}).$
with $\lambda_{\delta g}$ defined by (20). This expression for $\Xi$ arises as the degree $1$ term in $\epsilon$ in $\check{Y}^{\mathrm{glob},(1)}$, see (29).
@@
(15) $\check{W}=\langle[\delta A],\mathsf{a}^{2}\rangle+\frac{\epsilon}{2}\langle\mathsf{a}^{2},\check{\Theta}\mathsf{a}^{2}\rangle+\sum_{\Gamma}\frac{\epsilon^{l(\Gamma)-1}}{|\mathrm{Aut}(\Gamma)|}\check{\Phi}_{\Gamma}\quad\in\Omega^{\bullet}(U,\widehat{\mathrm{Sym}}(H_{A}^{\bullet}[1])^{*})[[\epsilon]]$
@@
(16) $\check{K}=K-\epsilon KHK+\epsilon^{2}KHKHK\in\bigoplus_{p=0}^{2}\Omega^{p}(U,\mathrm{End}(\Omega^{\bullet}(M,\mathfrak{g}))_{-1-p})[\epsilon]$
@@
(17) $\check{i}(\mathsf{a})=i(\mathsf{a})-\epsilon KHi(\mathsf{a})+\epsilon^{2}KHKHi(\mathsf{a})+\epsilon K\delta A\in\Omega^{\bullet+1}(U\times M,\mathfrak{g})[\epsilon].$
@@
(18) $\check{\Theta}=pHKHi\in\Omega^{2}(U,\mathrm{End}(H^{\bullet}_{A})_{-1}).$
@@
Here the first term in the brackets is the Grothendieck connection $\nabla^{G}=\nabla^{\mathbb{H}}+\{R_{\delta A},\cdot\}+\Delta_{\mathsf{a}}R_{\delta A}$ associated to the natural sum-over-trees formal exponential map [^14] See Appendix A.1 for a definition of formal exponential maps and their associated Grothendieck connections, and Appendix A.2.1 for the definition of the sum-over-trees exponential map on $\mathcal{M}^{\prime}$. on $\mathcal{M}^{\prime}$ (Lemma A.1). The 1-form $\alpha$ is obtained by restriction of $\{R_{\delta g},\cdot\}+\Delta_{\mathsf{a}}R_{\delta g}$, making the expression in brackets a flat superconnection. [^15] In fact, a restriction of the Grothendieck connection on $\mathcal{M}^{\prime}\times\mathrm{Met}$, plus $\Delta_{\mathsf{a}}$ term, see Lemma A.2. Finally, the remaining terms in (24) cancel when passing to the moduli space. [^16] There is a subtlety in this proof having to do with the restriction to harmonic tangent vectors (and the fact that the notion of a harmonic form changes when one moves on the moduli space), for which we refer to [8, Lemma 5.11].
@@
Expanding in form degrees along $\mathrm{Met}$, we have $\check{Y}^{\mathrm{glob}}=\check{Y}^{\mathrm{glob},(0)}+\check{Y}^{\mathrm{glob},(1)}+\cdots$ with $\check{Y}^{\mathrm{glob},(0)}=e^{c(\epsilon)S_{\mathrm{grav}}(g,\phi)}Y$ – the integrand in (10). [^18] Formula (8) for $Y$ was obtained in [8] from the computation of the r.h.s. of (28) where the key ingredient is the explicit formula for the projection (27) coming from homological perturbation lemma: $\mathsf{p}(\cdots)=e^{\langle\frac{\partial}{\partial\mathsf{a}^{2}},\frac{\partial}{\partial[\delta A]}\rangle}(\cdots)|_{\mathsf{a}^{1}=0}$. The exponential of the differential operator in the r.h.s. can also be written as the Berezin integral (8). We exploit the fact that one can use the odd Fourier transform to identify half-densities on $T^{*}[-1]\mathcal{M}^{\prime}$ equipped with BV differential $\Delta_{\mathcal{M}^{\prime}}$ with forms on $\mathcal{M}^{\prime}$ equipped with de Rham differential $d_{\mathcal{M}^{\prime}}$, see (30) below. [^19] See [14], [10], [11] for details.

### v2

@@ 1. Introduction
In his pioneering work [14] Witten was studying the path integral of Chern-Simons theory
@@ 2. The construction of the invariant
Let $M$ be a closed oriented smooth 3-manifold. Fix $G$ a compact simply-connected Lie group with Lie algebra $\mathfrak{g}$. We fix some invariant nondegenerate bilinear form $\langle,\rangle$ on $\mathfrak{g}$. [^2] E.g., the trace form associated to some representation. For $G$ simple, it is conventional in the context of Chern-Simons theory to choose $\langle,\rangle$ to be $\frac{1}{h^{\vee}}$ times the Killing form, with $h^{\vee}$ the dual Coxeter number. In this case, $\frac{1}{8\pi^{2}}\langle x,x\rangle$ corresponds under Chern-Weil homomorphism to a generator of $H^{4}(BG,\mathbb{Z})$, which is related to invariance of $e^{\frac{ik}{2\pi}S_{CS}(A)}$ under large gauge transformations for $k\in\mathbb{Z}$, cf. [14], [2]. Also fix a Riemannian metric $g$ on $M$ (this is auxiliary data: the invariants we construct will turn out to be independent of $g$).
@@
(5) $\Phi_{\Gamma}=\\ (-1)^{\mu}\int_{C_{V}(M)}\Big\langle\bigwedge_{\mathrm{edges\;}e=(uv)}\pi_{uv}^{*}\widehat{\eta}\;\wedge\;\bigwedge_{\mathrm{short\;loops\;}e=(vv)}\pi^{*}_{v}\widehat{\eta}^{\Delta}\;\wedge\;\bigwedge_{\mathrm{leaves\;}l}\pi^{*}_{v(l)}\widehat{i}(\mathsf{a}^{2}),\bigotimes_{\mathrm{vertices}}f\Big\rangle_{\Gamma}.$
@@
  • $\langle-,-\rangle_{\Gamma}$ is the extension of the pairing on $\mathfrak{g}$ to the pairing between $\mathfrak{g}^{\otimes N}$ and $\mathfrak{g}^{\otimes N}$ with $N$ the number of half-edges in $\Gamma$ according to the incidence of half-edges to vertices in $\Gamma$. Put another way, $\langle-,-\rangle_{\Gamma}=\langle\sigma_{\Gamma}(-),$-$\rangle$, where the pairing on the right matches $i$-th copy of $\mathfrak{g}$ to $i$-th copy of $\mathfrak{g}$ and $\sigma_{\Gamma}\in S_{N}$ is the permutation realizing the graph, see e.g. [8, Section 3.10.2]. [^7] One can think of a trivalent graph with leaves as a double coset $\Gamma\in{(S_{V}\ltimes S_{3}^{V})\backslash S_{N}/((S_{E}\ltimes\mathbb{Z}_{2}^{E})\times S_{\#\mathrm{leaves}})}$. In particular, a graph can be thought of as an equivalence class of permutations $\sigma_{\Gamma}$.
  • $(-1)^{\mu}$ is a sign depending on the graph $\Gamma$:
    (8) $(-1)^{\mu}=(-1)^{E}(-1)^{\left[\frac{V}{2}\right]}(-1)^{\sigma_{\Gamma}}$
    where the last factor is the sign of the permutation $\sigma_{\Gamma}$ above, $E$ is the number of edges. [^8] A comment on the origin of the factors in the sign: $(-1)^{E}$ corresponds to the fact that if we are weighing the graph with $\epsilon^{-\chi(\Gamma)}$, with $\chi$ the Euler characterisitic, then the edge should be decorated with $\widehat{\eta}$ rather than $\widehat{\eta}$ (or equivalently: vertices are decorated with $\frac{1}{\epsilon}f$ and edges with $\epsilon\widehat{\eta}$). $(-1)^{\sigma_{\Gamma}}$ is as in (3.10.2) in [8]. The middle factor is a Koszul sign arising from reshuffling of $V$ integrals over $M$ into a single integral over $M^{\times V}$ (which can be replaced – by cutting out diagonals and compactifying – by $C_{V}(M)$) in the free theory correlator (Gaussian average) $\langle\int_{M}\langle A,[A,A]\rangle\cdots\int_{M}\langle A,[A,A]\rangle\rangle_{0}$.
@@
$\Theta=-\frac{1}{2\epsilon}\int_{M}\langle\widehat{i}(\mathsf{a}^{2})\stackrel{{\scriptstyle\wedge}}{{,}}d\widehat{i}(\mathsf{a}^{2})\rangle=-\frac{\epsilon}{2}\langle\mathsf{a}^{2},p\,\mathrm{ad}^{*}_{i(\zeta)}GdG\mathrm{ad}^{*}_{i(\zeta)}i(\mathsf{a}^{2})\rangle,$
@@
We define the “Chern-Simons volume form” on $\mathcal{M}^{\prime}$ as the Berezin integral [^10] We use the following sign convention for the Berezinian in (9): $\mathcal{D}\mathsf{a}^{2}\,\mathcal{D}\zeta=\prod_{i}\mathcal{D}\mathsf{a}^{2}_{i}\,\mathcal{D}\zeta^{i}$ where $\zeta^{i}$ are coordinates on $H^{1}_{A}[1]$ associated to some basis $\xi_{i}$ in $H^{1}_{A}$ and $\mathsf{a}^{2}_{i}$ are coordinates on $H^{2}_{A}[-1]$ associated to the dual basis $\xi^{i}$ in $H^{2}_{A}$. In our notations, for an odd line $\Pi\mathbb{R}$ parametrized by odd coordinate $\psi$, $\mathcal{D}\psi$ is the standard coordinate Berezinian satisfying $\int_{\Pi\mathbb{R}}\mathcal{D}\psi\;\psi=1$.
(9) $Y=\tau_{A}^{\frac{1}{2}}\int_{H_{A}^{2}[-1]\oplus H_{A}^{1}[1]}\mathcal{D}\mathsf{a}^{2}\,\mathcal{D}\zeta\;e^{\langle\zeta,\mathsf{a}^{2}\rangle+W-W_{0}}\quad\in\Omega^{\mathrm{top}}(\mathcal{M}^{\prime})[[\epsilon]].$
@@
be the Chern-Simons functional evaluated on the Levi-Civita connection associated to the metric $g$, trivialized using $\phi$. Here $\mathrm{tr}$ is normalized as minus half the trace of adjoint representation of $\mathfrak{so}(3)$.
@@
(10) $c(\epsilon)=-\frac{\langle f,f\rangle}{24(2\pi)^{2}}\epsilon+O(\epsilon^{3})\;\in\epsilon\mathbb{R}[[\epsilon^{2}]]$
@@
does not depend on the metric $g$ and thus is an invariant of the framed oriented 3-manifold $(M,\phi)$. In (10), $\langle f,f\rangle$ is the value of the $\Theta$-graph as a Jacobi graph for $\mathfrak{g}$. [^12] If one uses the pairing on $\mathfrak{g}$ given by $\frac{1}{h^{\vee}}$ times the Killing form (cf. footnote 2), then $\langle f,f\rangle=-h^{\vee}\dim\mathfrak{g}$.
@@ 3. The lowest (nontrivial) degree invariant
(14) $\Lambda_{1}=\int_{\mathcal{M}^{\prime}_{\alpha}}\tau_{A}^{\frac{1}{2}}\left(\Phi_{1}+\Phi_{2}+\Phi_{3}+\Phi_{4}+\Phi_{5}+\Phi_{6}-\frac{\langle f,f\rangle}{24(2\pi)^{2}}S_{\mathrm{grav}}(g,\phi)\right)$
@@
$\Phi_{1}$ $=$ $-\frac{1}{12}\int_{C_{2}(M)}f^{abc}f^{a^{\prime}b^{\prime}c^{\prime}}\eta^{aa^{\prime}}\wedge\eta^{bb^{\prime}}\wedge\eta^{cc^{\prime}},$
$\Phi_{2}$ $=$ $\frac{1}{8}\int_{C_{2}(M)}f^{abc}f^{a^{\prime}b^{\prime}c^{\prime}}\pi_{1}^{*}(\eta^{\Delta})^{ab}\wedge\eta^{cc^{\prime}}\wedge\pi_{2}^{*}(\eta^{\Delta})^{a^{\prime}b^{\prime}},$
$\Phi_{3}$ $=$ $-\frac{1}{2}\int_{M}f^{abc}([K\mathrm{ad}^{*}_{\chi_{i}}G]^{\Delta})^{ab}\wedge\chi^{ic},$
$\Phi_{4}$ $=$ $\frac{1}{2}\int_{M}f^{abc}(\eta^{\Delta})^{ab}\wedge(G\mathrm{ad}^{*}_{\chi_{i}}\chi^{i})^{c},$
$\Phi_{5}$ $=$ $\frac{1}{2}\int_{M}\langle\chi^{i}\stackrel{{\scriptstyle\wedge}}{{,}}\mathrm{ad}^{*}_{\chi_{j}}GdG\mathrm{ad}^{*}_{\chi_{i}}\chi^{j}\rangle,$
$\Phi_{6}$ $=$ $-\frac{1}{2}\int_{M}\langle\chi^{i}\stackrel{{\scriptstyle\wedge}}{{,}}\mathrm{ad}^{*}_{\chi_{i}}GdG\mathrm{ad}^{*}_{\chi_{j}}\chi^{j}\rangle.$
Indices $a,b,c,\ldots$ refer to an orthonormal basis in $\mathfrak{g}$; summation over repeated indices is implied. [^14] Alternatively one can use an orthonormal basis with respect to minus the pairing on $\mathfrak{g}$, which is convenient when the pairing is negative-definite. Then one should flip the signs of $\Phi_{3},\Phi_{4}$. For each point of $\mathcal{M}^{\prime}_{\alpha}$ we choose arbitrarily a basis $\chi_{i}$ in $A$-harmonic 1-forms, and $\chi^{i}$ is the dual basis of harmonic 2-forms. The object $K\mathrm{ad}^{*}_{\chi_{i}}G]^{\Delta}\in\Omega^{1}(M,\mathfrak{g}\otimes\mathfrak{g})$ is the (regularized) evaluation of the integral kernel of the operator in brackets on the diagonal. [^15] One can also write $\Phi_{3}$ as an integral over the two-point configuration space $\frac{1}{2}\int_{C_{2}(M)}f^{abc}\eta^{da}\wedge[\mathrm{ad}^{*}_{\chi_{i}}G]^{db}\wedge\pi_{2}^{*}\chi^{ic}$, with $\cdots]\in\Omega^{2}(C_{2}(M),\mathfrak{g}\otimes\mathfrak{g})$ standing for the integral kernel of the operator in brackets. Yet another way to write $\Phi_{3}$ is as a regularized supertrace $\frac{1}{2}\mathrm{Str}_{\Omega^{\bullet}(M,\mathfrak{g})}^{\mathrm{reg}}\left(K\mathrm{ad}^{*}_{\chi_{i}}G\mathrm{ad}_{\chi^{i}}\right)$. Expressions $\Phi_{1},\ldots,\Phi_{6}$ correspond to the six graphs in Figure 1. Graphs $\Gamma_{1},\Gamma_{2}$ are as in the usual Axelrod-Singer two-loop Chern-Simons invariant (at an acyclic flat connection), whereas the remaining graphs $\Gamma_{3},\ldots,\Gamma_{6}$ account (in an appropriate sense) for the effect of globalization on the moduli space $\mathcal{M}^{\prime}$.
@@
(15) $\Xi=\iota_{\chi_{i}}\tau_{A}^{\frac{1}{2}}\cdot\Big(\int_{M}\frac{1}{2}f^{abc}(\eta^{\Delta})^{ab}\wedge(K\lambda_{\delta g}\chi^{i})^{c}+\int_{M}\frac{1}{2}f^{abc}([K\lambda_{\delta g}K]^{\Delta})^{ab}\wedge\chi^{ic}\\ +\int_{M}\langle\chi^{j}\stackrel{{\scriptstyle\wedge}}{{,}}\lambda_{\delta g}G\mathrm{ad}^{*}_{\chi_{j}}\chi^{i}\rangle-\int_{M}\langle\chi^{j}\stackrel{{\scriptstyle\wedge}}{{,}}\mathrm{ad}^{*}_{\chi_{j}}G\lambda_{\delta g}\chi^{i}\rangle\Big)\quad\in\Omega^{\mathrm{top}-1}(\mathcal{M}^{\prime}_{\alpha}).$
[…]
@@
(16) $\check{W}=-\langle[\delta A],\mathsf{a}^{2}\rangle+\frac{\epsilon}{2}\langle\mathsf{a}^{2},\check{\Theta}\mathsf{a}^{2}\rangle+\sum_{\Gamma}\frac{\epsilon^{l(\Gamma)-1}}{|\mathrm{Aut}(\Gamma)|}\check{\Phi}_{\Gamma}\quad\in\Omega^{\bullet}(U,\widehat{\mathrm{Sym}}(H_{A}^{\bullet}[1])^{*})[[\epsilon]]$
@@
(17) $\check{K}=K+\epsilon KHK+\epsilon^{2}KHKHK\in\bigoplus_{p=0}^{2}\Omega^{p}(U,\mathrm{End}(\Omega^{\bullet}(M,\mathfrak{g}))_{-1-p})[\epsilon]$
@@
(18) $\check{i}(\mathsf{a})=i(\mathsf{a})+\epsilon KHi(\mathsf{a})+\epsilon^{2}KHKHi(\mathsf{a})+\epsilon K\delta A\in\Omega^{\bullet+1}(U\times M,\mathfrak{g})[\epsilon].$
@@
(19) $\check{\Theta}=-pHKHi\in\Omega^{2}(U,\mathrm{End}(H^{\bullet}_{A})_{-1}).$
@@
Here the first term in the brackets is the Grothendieck connection $\nabla^{G}=\nabla^{\mathbb{H}}+\{R_{\delta A},\cdot\}+\Delta_{\mathsf{a}}R_{\delta A}$ associated to the natural sum-over-trees formal exponential map [^20] See Appendix A.1 for a definition of formal exponential maps and their associated Grothendieck connections, and Appendix A.2.1 for the definition of the sum-over-trees exponential map on $\mathcal{M}^{\prime}$. on $\mathcal{M}^{\prime}$ (Lemma A.1). The 1-form $\alpha$ is obtained by restriction of $\{R_{\delta g},\cdot\}+\Delta_{\mathsf{a}}R_{\delta g}$, making the expression in brackets a flat superconnection. [^21] In fact, a restriction of the Grothendieck connection on $\mathcal{M}^{\prime}\times\mathrm{Met}$, plus $\Delta_{\mathsf{a}}$ term, see Lemma A.2. Finally, the remaining terms in (25) cancel when passing to the moduli space. [^22] There is a subtlety in this proof having to do with the restriction to harmonic tangent vectors (and the fact that the notion of a harmonic form changes when one moves on the moduli space), for which we refer to [9, Lemma 5.11].
@@
Expanding in form degrees along $\mathrm{Met}$, we have $\check{Y}^{\mathrm{glob}}=\check{Y}^{\mathrm{glob},(0)}+\check{Y}^{\mathrm{glob},(1)}+\cdots$ with $\check{Y}^{\mathrm{glob},(0)}=e^{c(\epsilon)S_{\mathrm{grav}}(g,\phi)}Y$ – the integrand in (11). [^24] Formula (9) for $Y$ was obtained in [9] from the computation of the r.h.s. of (29) where the key ingredient is the explicit formula for the projection (28) coming from homological perturbation lemma: $\mathsf{p}(\cdots)=\Big(e^{\langle\frac{\partial}{\partial\mathsf{a}^{2}},\frac{\partial}{\partial[\delta A]}\rangle}(\cdots)\Big)\Big|_{\mathsf{a}^{1}=0,\,[\delta A]=0}$. The exponential of the differential operator in the r.h.s. can also be written as the Berezin integral (9). We exploit the fact that one can use the odd Fourier transform to identify half-densities on $T^{*}[-1]\mathcal{M}^{\prime}$ equipped with BV differential $\Delta_{\mathcal{M}^{\prime}}$ with forms on $\mathcal{M}^{\prime}$ equipped with de Rham differential $d_{\mathcal{M}^{\prime}}$, see (31) below. [^25] See [15], [11], [12] for details.